# Comparison of bar formation mechanisms II: does a tidally induced bar grow faster than an internally developed bar?

### v1

@@ Appendix A Fitting validation
In addition to Equation 3, we explored alternative exponential formulas to fit the $A_{2}$ evolution. One such formula is:
@@
which offers greater flexibility in the fitting process. We also examined the following formula:
@@
This choice is driven by the consideration that the bar strength $A_{2}$ should be zero at $t=0$, as the disks are initialized to be axisymmetric at the beginning of the simulations.
@@
To further demonstrate the reliability of the exponential fitting, we compare the fitted $t_{0}$ values with the true $t_{0}$ values in Figure 9. The fitted $t_{0}$ is defined as the time at which the fitted $A_{2}$ profile reaches 0.1, while the true $t_{0}$ is the time when the actual $A_{2}$ profile attains this value. The majority of the simulations show a nice agreement within ±$\pm 0.1$ Gyr between and fitted and true $t_{0}$. Only 6 simulations show slightly larger offsets, ranging between 0.1 and 0.3  Gyr. This agreement underscores the robustness of the exponential fitting approach.

### v3

@@ Appendix A Fitting validation
In addition to Equation 4, we explored alternative exponential formulas to fit the $A_{2}$ evolution. One such formula is
@@
which offers greater flexibility in the fitting process. We also examined the following formula
@@
This choice is driven by the consideration that the bar strength $A_{2}\approx 0$ at $t=0$ if the disk is set to be perfectly axisymmetric at the beginning of the simulations, even though $A_{2}$ is not expected to stay at zero for a long time.
@@
To further demonstrate the reliability of the exponential fitting, we compare the fitted $t_{0}$ values with the true $t_{0}$ values in Figure 9. The fitted $t_{0}$ is defined as the time at which the fitted $A_{2}$ profile reaches 0.1, while the true $t_{0}$ is the time when the actual $A_{2}$ profile attains this value. The majority of the simulations show good agreement within ±$\pm 0.1$ Gyr between and fitted and true $t_{0}$. Only six simulations show slightly larger offsets, ranging between 0.1 and 0.3  Gyr. This agreement underscores the robustness of the exponential fitting approach.